The reaction H2(g) + I2(g) <=> 2HI(g) has Kc = 50.2 at 718K. If a flask is charged with 4.25M HI, what are the equilibrium concentrations of H2 and I2? at equilibrium, [H^2}+[I^2]= ?????
Answer #1

ICE Table:

Equilibrium constant expression is

Kc = [HI]^2/[H2]*[I2]

50.2 = (4.25-2*x)^2/(1*x)^2

sqrt(50.2) = (4.25-2*x)/(1*x)

7.085 = (4.25-2*x)/(1*x)

7.085*x = 4.25-2*x

-4.25 + 9.085*x = 0

x = 0.468

At equilibrium:

[H2] = x = 0.468 M

[I2] = x = 0.468 M

Answer:

[H2] = 0.468 M

[I2] = 0.468 M